# Clothing agnostic Pre-inpainting Virtual Try-ON

### v1

@@ Abstract
With the development of deep learning technology, virtual try-on technology has become an important application value in the fields of e-commerce, fashion, and entertainment. The recently proposed Leffa has improved the texture distortion problem of diffu-sion-based models, but there are limitations in that the bottom detection inaccuracy and the existing clothing silhouette remain in the synthesis results. To solve this problem, this study proposes CaP-VTON (Clothing agnostic Pre-inpainting Virtual Try-ON). CaP-VTON has improved the naturalness and consistency of whole-body clothing syn-thesis by integrating multi-category masking based on Dress Code and skin inpainting based on Stable Diffusion. In particular, a generate skin module was introduced to solve the skin restoration problem that occurs when long-sleeved images are converted into short-sleeved or sleeveless ones, and high-quality restoration was implemented consider-ing the human body posture and color. As a result, CaP-VTON recorded 92.5%, which is 15.4% better than Leffa in short-sleeved synthesis accuracy, and showed the performance of consistently reproducing the style and shape of reference clothing in visual evaluation. These structures maintain model-agnostic properties and are applicable to various diffu-sion-based virtual inspection systems, and can contribute to applications that require high-precision virtual wearing, such as e-commerce, custom styling, and avatar creation.
@@ 1 Introduction
However, there were synthesis results that could not be covered by the existing Leffa model, and we supplemented them by adding several things from the pipelines of the ex-isting study. The points that can be supplemented by the existing Leffa model are shown in Figure 1. The first can be confirmed in Figure 1(a). Since the existing Leffa model is mainly optimized for virtual inspection centered on upper-body clothes, detection and masking of the lower-body areas are relatively vulnerable. This can put great restrictions on this part in a virtual inspection scenario centered on whole-body clothing or bottoms. In particular, when the lower-body could not be detected in the input image, there was a limitation in that the whole body was synthesized even if the upper-body clothes were synthesized. Therefore, a separate specialized preprocessing strategy that elaborately re-flects this is needed. The second was that the existing clothing characteristics included in the input source image had more than a certain level of influence on the final output result. Figure 1(b) was outputted as a Leffa-Corrective Masking version that improved the first limitation to confirm independently of the first problem. When trying to synthesize short-sleeved clothing with the image of a person wearing a long-sleeved top as shown in Figure 1(b), it was observed that the existing long-sleeved shape was maintained regard-less of the input garment. This seemed to be because the skin area to be exposed, such as the forearm, shoulder, and neck, was not sufficiently restored in the process of removing clothing, and the existing clothing type acted as an interference factor in the subsequent clothing synthesis process. This phenomenon could lead to unnatural results in the syn-thesis between clothing with large differences in shape such as sleeve length and neckline.
@@
By integrating the Dress Code-based improved masking mechanism and skin inpainting pipeline while maintaining the advantages of the existing Leffa pipeline, this study aims to achieve complete synthesis results in all clothing categories and bottoms. This, combined with Leffa’s excellent attention-based flow field learning ability, was able to present an integrated solution that satisfies the three goals at the same time: detailed texture preservation, full-body clothing synthesis, and input image independent clothing replacement. In particular, it was confirmed that the interference phenomenon caused by existing clothing characteristics was significantly improved along with the morphological inconsistency, the unnaturalness of the boundary, and the incongruity with the body ratio shown in the bottom synthesis. In addition, this study is of high practical value in that it provides a universal improvement method that can be applied to various diffusion-based models through improved accuracy of the masking process and improved skin restoration quality while maintaining the model-agnostic characteristics of the existing Leffa model. This approach is differentiated from existing studies that have presented solutions de-pendent on specific models or datasets, and is expected to be widely used in various ap-plications in the virtual try-on field in the future.
@@ 2.2 Virtual Try-on Models
Later, the introduction of transformers and attention mechanisms enabled more so-phisticated feature matching and spatial relationship modeling even in Virtual Try-On. The multi-head self-attention proposed in Vaswani, A. et al.[14] became the basis of the transformer model, and models that use it in Virtual Try-On also began to emerge. Cat-VTON[15] selectively fine-tuned only the self-attention module within the diffusion U-Net and conducted lightweight learning with 49.57M parameters, removing text encoders and cross-attention, and maximizing efficiency by simply concatenating and inputting human and costume images in the spatial dimension. However, detailed semantic alignment control was limited. Leffa[1] improved the self-attention and cross-attention map by ap-plying flow field regularization to the attachment of the diffusion model. This enabled more sophisticated attention control than the existing mask inpainting method, zero cross-attention method, and prompt-aware mask method, but there were still problems with the complexity and fine-detail distortion of flow field regularization.
@@ 3.2 Stable Diffusion Generate Skin Inpainting: Removing existing clothing silhouette
When trying to synthesize short-sleeved clothing with the image of a person wearing a long-sleeved top in the existing model, there was a problem that the existing long-sleeved shape was maintained regardless of the type of reference clothing as shown in Figures 3(A) and (C), and the problem that the silhouette of the existing clothing re-mained as shown in Figure 3(B). This was because the silhouette of the existing clothing interfered with the synthesis of new clothing because the exposed skin areas such as the forearms, shoulders, and neck were not properly restored during the clothing removal process.
In order to essentially solve this problem, the generate skin method pipeline has adopted an approach that inserts short-sleeved clothes and realistic skin into the area where existing clothes have been removed, which does not affect the output. This was in-tended to accurately reflect the unique shape and style of the reference clothes regardless of the existing clothing characteristics of the input image.
@@ 3.2.1 Generating input information to build synthetic model
In order to perform inpainting through the Generate Skin function, an OpenPose im-age containing the inpainting mask image and the pose information of the input image is required. The Generate Skin Inpainting Mask Output of Figure 4 is generated using the masking pipeline provided by Leffa’s DressCode model. In addition, the OpenPose Output of Figure 5 is used as conditioning information of the ControlNet in the process of Stable Diffusion-based inpainting. As shown in Figure 5, when the ControlNet is not used, the pose identity of the original person is not accurately transmitted to the network, resulting in a problem of creating a structurally distorted image, such as a misalignment of the body position such as the arm or shoulder or mixing with the background. As a result, we determined that the problem of preserving the pose and securing the accuracy of the skin area in the process of inpainting directly affects the realism of the final result, and to solve this problem, the Generate Skin Inpainting Mask and OpenPose-based ControlNet[18] in-formation were used as input information for the Generate Skin function. In addition, as can be seen in Figure 5, there were cases where the color of the inpainted skin did not properly reflect the skin tone of the original image, resulting in the problem of creating darker or brighter skin than it actually is. To compensate for this, the Skin Tone Extraction technique was introduced to create a skin texture that naturally harmonizes with the original image. This technique is implemented by detecting the skin area in the YCrCb color space and calculating the average skin tone value in the HSV color space. The skin area created through this is naturally blended with the skin tone of the original image, so that the overall visual consistency of the resulting image can be maintained.
@@ 3.3 Modified Leffa pipeline: Multi-category enabled based on DressCode
Figure 7: Modified Ensemble model of VITON-HD and DressCode Model
@@ 3.3.1 Characteristic of VITON-HD and DressCode Model
Table 1: Numerical Comparison between two models.
@@ 4.1 Experimental Environment
This study was experimented in the Google Colab environment, and Ubuntu 20.04 LTS was used as an operating system. The programming language consisted of Python 3.11.13 version, and the compiler was GCC 11.4.0. As the main library, packages related to image processing and deep learning such as OpenCV and TorchGeometry were used, centering on PyTorch (version 1.6.0 and higher, GPU acceleration support). Stable Diffu-sion was used as a model for image generation, and the image was created in a 30-step setting based on the Diffuser library or similar framework. The experimental environment was operated through the Colab interface based on Jupiter Notebook. The hardware con-figuration consisted of NVIDIA A100 GPU (40GB memory), 83.5GB of system RAM, and Intel Xeon-based CPU, and the disk space used in the experiment was about 80GB. When Stable Diffusion was configured to a 30-step setting, it took about 8 seconds on average to generate a single image in the NVIDIA A100 GPU environment. The data using the model production and evaluation were made using the Viton-HD dataset, 1100 clothing images, and 1100 people images. The person used to create the resultant image was produced through a generative model.
@@ 4.2.1 Visual Comparison of the synthesis results
4.2.2 Quantitative Comparison
Table 2 is the result of comparing the performance between the existing virtual wear-ing models and the proposed CaP-VTON model with FID, SSIM, and LPIPS indicators based on the category of the award. The proposed CaP-VTON model showed a slightly lower quantitative performance with FID 11.46, SSIM 0.8573, and LPIPS 0.0849 due to the additional image inference pipeline(Generate Skin Method).
However, these figures are general indicators focusing on overall image quality and perception-based similarity, and there is a limit to evaluating the accuracy of conversion from long-sleeve to short-sleeve, which is the purpose of this study. Therefore, observa-tion-based accuracy was added and evaluated.
@@ 4.2.2 Quantitative Comparison
Figure 11: Comparison the percentage of correct ouput
4.2.3 Evaluation based on Short-Sleeve Synthesis Probability
This study additionally conducted a quantitative experiment to evaluate whether or not the reference clothing was actually synthesized as a normal short sleeve when it was short-sleeved, targeting the long-sleeved wearing input image. This is to quantitatively verify the form-consistency of the model’s clothing shape by checking how much the sil-houette of the existing clothing interferes with the final output. As shown in Figure 10, it can be seen that the Leffa model is very vulnerable to this part. Figure 11 is a visualization of the results of classifying the output results into normal output (when a short-sleeved silhouette matching the reference clothing is accurately generated) and abnormal output (when a long-sleeved silhouette remains or short-sleeved clothing is synthesized as long-sleeved clothing).
@@ 4.2.3 Evaluation based on Short-Sleeve Synthesis Probability
This result means that CaP-VTON recorded about 15.4%p higher accuracy of clothing silhouette consistency than Leffa. In particular, CaP-VTON showed excellent performance in the ability to accurately reflect short-sleeved reference clothing after removing long sleeves and naturally restoring the skin of the arms and shoulders, which can be said to be the result of empirically proving the effectiveness of generate skin-based pre-processing and improved masking pipeline.
@@
On the other hand, Leffa showed excellent quantitative performance in terms of FID and SSIM, but the above experimental results show the limitation that it is difficult to fully explain the suitability of the model’s clothing replacement only with quantitative image quality evaluation. In actual applications, the consistency-based evaluation method pro-posed in this study is a very important measure in a practical sense because the accuracy and visual naturalness of the clothing shape conversion desired by the user act as more important factors.
@@
In conclusion, CaP-VTON showed a distinct improvement over Leffa in terms of structural consistency, meaning that the pipeline was designed to minimize dependence on existing clothing types and accurately reflect the style of reference clothing. These re-sults support the feasibility of CaP-VTON in high-dimensional applications such as full-body virtual wear and style transformation.

### v3

@@ Abstract
With the development of deep learning technology, virtual try-on technology has devel-oped important application value in the fields of e-commerce, fashion, and entertainment. The recently proposed Leffa technology has addressed the texture distortion problem of diffusion-based models, but there are limitations in that the bottom detection inaccuracy and the existing clothing silhouette persist in the synthesis results. To solve this problem, this study proposes CaP-VTON (Clothing Agnostic Pre-Inpainting Virtual Try-On). CaP-VTON integrates DressCode-based multi-category masking and Stable Diffu-sion-based skin inflation preprocessing; in particular, a generated skin module was in-troduced to solve skin restoration problems that occur when long-sleeved images are con-verted to short-sleeved or sleeveless ones, introducing a preprocessing structure that im-proves the naturalness and consistency of full-body clothing synthesis, and allowing the implementation of high-quality restoration considering human posture and color. As a result, CaP-VTON achieved 92.5%, which is 15.4% better than Leffa, in short-sleeved syn-thesis accuracy, and consistently reproduced the style and shape of the reference clothing in visual evaluation. These structures maintain model-agnostic properties and are appli-cable to various diffusion-based virtual inspection systems; they can also contribute to applications that require high-precision virtual wearing, such as e-commerce, custom styling, and avatar creation.
@@ 1 Introduction
However, there were synthesis results that could not be covered by the existing Leffa model, and we supplemented them by adding several things from the pipelines of the existing study. The points that can be supplemented by the existing Leffa model are shown in Figure 1. The first can be confirmed in Figure 1(a). Since the existing Leffa model is mainly optimized for virtual inspection centered on upper-body clothes, detection and masking of the lower-body areas are relatively vulnerable. This can put great restrictions on this part in a virtual inspection scenario centered on whole-body clothing or bottoms. In particular, when the lower-body could not be detected in the input image, there was a limitation in that the whole body was synthesized even if the upper-body clothes were synthesized. Therefore, a separate specialized preprocessing strategy that elaborately reflects this is needed. The second was that the existing clothing characteristics included in the input source image had more than a certain level of influence on the final output result. Figure 1(b) was outputted as a Leffa-Corrective Masking version that improved the first limitation to confirm independently of the first problem. When trying to synthesize short-sleeved clothing with the image of a person wearing a long-sleeved top as shown in Figure 1(b), it was observed that the existing long-sleeved shape was maintained regardless of the input garment. This seemed to be because the skin area to be exposed, such as the forearm, shoulder, and neck, was not sufficiently restored in the process of removing clothing, and the existing clothing type acted as an interference factor in the subsequent clothing synthesis process. This phenomenon could lead to unnatural results in the synthesis between clothing with large differences in shape such as sleeve length and neckline.
@@
By integrating the Dress Code-based improved masking mechanism and skin inpainting pipeline while maintaining the advantages of the existing Leffa pipeline, this study aims to achieve complete synthesis results in all clothing categories and bottoms. This, combined with Leffa’s excellent attention-based flow field learning ability, was able to present an integrated solution that satisfies the three goals at the same time: detailed texture preservation, full-body clothing synthesis, and input image independent clothing replacement. In particular, it was confirmed that the interference phenomenon caused by existing clothing characteristics was significantly improved along with the morphological inconsistency, the unnaturalness of the boundary, and the incongruity with the body ratio shown in the bottom synthesis. In addition, this study is of high practical value in that it provides a universal improvement method that can be applied to various diffusion-based models through improved accuracy of the masking process and improved skin restoration quality while maintaining the model-agnostic characteristics of the existing Leffa model. This approach is differentiated from existing studies that have presented solutions dependent on specific models or datasets, and is expected to be widely used in various applications in the virtual try-on field in the future.
@@ 2.2 Virtual Try-on Models
Later, the introduction of transformers and attention mechanisms enabled more so-phisticated feature matching and spatial relationship modeling even in Virtual Try-On. The multi-head self-attention proposed in Vaswani, A. et al.[14] became the basis of the transformer model, and models that use it in Virtual Try-On also began to emerge. Cat-VTON[15] selectively fine-tuned only the self-attention module within the diffusion U-Net and conducted lightweight learning with 49.57M parameters, removing text encoders and cross-attention, and maximizing efficiency by simply concatenating and inputting human and costume images in the spatial dimension. However, detailed semantic alignment control was limited.
TryOnDiffusion[16] achieved precise garment-person matching by leveraging garment-aware person representations and cross-attention based on latent diffusion. It recorded an FID of 23.3 on the VITON-HD dataset, demonstrating improved performance over existing diffusion models in terms of detailed clothing texture reproduction and pose preservation. HumanDiffusion[17] proposed a diffusion pipeline that separates human image representations into a distinct latent space and combines cross-attention with mask inpainting. It delivered natural results for full-body synthesis and multi-clothing category replacement, achieving FID scores ranging from 30.4 to 31.2 on the DeepFashion-Multimodal dataset.
Leffa[1] improved the self-attention and cross-attention map by ap-plying flow field regularization to the attachment of the diffusion model. This enabled more sophisticated attention control than the existing mask inpainting method, zero cross-attention method, and prompt-aware mask method, but there were still problems with the complexity and fine-detail distortion of flow field regularization.
@@ 3.2 Stable Diffusion Generate Skin Inpainting: Removing existing clothing silhouette
When trying to synthesize short-sleeved clothing from the image of a person wearing a long-sleeved top using the existing model, a problem arises in that the existing long-sleeved shape is maintained regardless of the type of reference clothing, as shown in Figure 3(A) and (C), and the silhouette of the existing clothing remains, as shown in Figure 3(B). This is because the silhouette of the existing clothing interferes with the synthesis of new clothing because the areas of exposed skin, such as the forearms, shoulders, and neck, are not properly restored during the clothing removal process.
To solve this problem, the generate skin method pipeline inserts short-sleeved clothes and realistic skin into the area where existing clothes have been removed, which does not affect the output. This is intended to accurately reflect the unique shape and style of the reference clothes regardless of the existing clothing characteristics of the input image.
The number of data points used in the garment silhouette removal experiment in the Generate Skin interposing module was set to 599 to simultaneously secure the reproduci-bility and the efficiency of the experiment in a limited GPU resource environment (Colab-based). “Wearing Held Tight Short Sleeve Shirt, high quality skin, realistic, high quality” was used for the prompt, and ”Blurry, low quality, artifacts, deformed, ugly, tex-ture, watermark, text, bad anatomy, extra limbs, face, hands, fingers” were used to pre-clude the creation of elements corresponding to the prompt. The inference step was set to 20 steps, and no separate fine-tuning was performed.
@@ 3.2.1 Generating input information to build synthetic model
To perform inpainting through the Generate Skin function, an OpenPose image con-taining the inpainting mask image and the pose information of the input image is re-quired. The Generate Skin Inpainting Mask Output of Figure 4 was generated using the masking pipeline provided by Leffa’s DressCode model. In addition, the OpenPose Output of Figure 5 has been used as conditioning information for the ControlNet in the process of Stable Diffusion-based inpainting. As shown in Figure 5, when the ControlNet is not used, the pose identity of the original person is not accurately transmitted to the network, re-sulting in the problem of creating a structurally distorted image, featuring, for example, a misalignment of the body position such as the arm or shoulder, or mixing with the back-ground. As a result, we determined that the problem of preserving the pose and securing the accuracy of the skin area in the process of inpainting directly affects the realism of the final result, and to solve this problem, information on the Generate Skin Inpainting Mask and OpenPose-based ControlNet [20] was used as the input information for the Generate Skin function. In addition, as can be seen in Figure 5, there were cases wherein the color of the inpainted skin did not properly reflect the skin tone of the original image, creating darker or brighter skin than was actually present. To compensate for this, the Skin Tone Extraction technique was introduced to create a skin texture that naturally harmonizes with that in the original image. This technique is implemented by detecting the skin area in the YCrCb color space and calculating the average skin tone value in the HSV color space. The skin area created through this approach is naturally blended with the skin tone of the original image, such that the overall visual consistency of the resulting image can be maintained.
@@ 3.2.2 MajicMix Realistic model based high quality skin synthesis
Figure 7: Modified Ensemble model of VITON-HD and DressCode Model
@@ 3.3.1 Characteristic of VITON-HD and DressCode Model
Table 1: Numerical Comparison between two models (paired : If the costume in the input person image and the costume image provided as a reference are the same, the result of the production can be compared directly to the correct image because the model actually produces the image in the same clothes. Therefore, we used quantitative image quality indicators such as SSIM and LPIPS in addition to the FID. unpaired : refers to the case where the costume in the input person image and the costume image given as reference are different. In this case, since there is no correct answer for the generated image, indicators such as SSIM or LPIPS cannot be used. Instead, we only used dis-tribution-based indicators such as FID, which can evaluate the overall quality of generation. In-dicators with down arrows have higher similarity with lower values, and indicators with up ar-rows have higher similarity with higher values.
@@ 4.1 Experimental Environment
This study was conducted in the Google Colab environment, and Ubuntu 20.04 LTS was used as an operating system. The programming language consisted of Python 3.11.13 version, and GCC 11.4.0 was used as the compiler. As the main library, packages related to image processing and deep learning such as OpenCV and TorchGeometry were used, centering on PyTorch (1.6.0 or higher, support for GPU acceleration).
As a generative model for image generation, the ChilloutMix checkpoint based on Stable Diffusion 3.5 was used. This is a derivative weight (checkpoint) that improves hu-man texture and lighting expression while maintaining the structure of Stable Diffusion 3.5, and was driven through the Diffusers library. The main settings in the generation process consisted of cfg scale = 7, steps = 25, sampler = Euler a, clip skip = 2. Person imag-es generated through this setting were used for matching and synthesis experiments with clothing images within the CaP-VTON pipeline.
The hardware configuration consisted of NVIDIA A100 GPU (40GB memory), 83.5GB system RAM, and Intel Xeon-based CPU, and the experiment was conducted through the Colab interface based on Jupyter Notebook. When the Stable Diffusion model was config-ured with the above settings, it took an average of about 8 seconds to generate a single image.
To assess whether the model generalizes effectively to new data and maintains stable performance on real human images, two types of person data were prepared. Specifically, 100 AI-generated person images were produced using the ChilloutMix model, and 1,000 real person images were obtained from the VITON-HD Test dataset. These were combined to create a unified person dataset containing 1,100 images, paired with a cloth dataset of equal size (1,100 garment images).
All images were normalized to a base resolution of 1024×768 to ensure reproducibility and computational efficiency under the limited GPU environment provided by Colab. As a result, the final experimental dataset consisted of 1,100 person images and 1,100 cloth images.
@@ 4.2.1 Visual Comparison of the synthesis results
4.2.2 Comparison for Image Synthesis Image Quantity
Table 2 compares performance between the existing virtual wearing models and the proposed CaP-VTON model, with FID, SSIM, and LPIPS indicators based on the assigned category. The proposed CaP-VTON model achieved a slightly lower quantitative performance, with FID 11.46, SSIM 0.8573, and LPIPS 0.0849, due to the addition of the image inference pipeline (Gen-erate Skin method).
When compared numerically only, our research shows moderate quality among the other synthetic models mentioned in the paper. This means that the synthesis result is not the best, and it does not produce heterogeneity to the degree to which humans perceive it. These figures are general indicators that focus on overall image quality and percep-tion-based similarity. In other words, there is a limit to evaluations of the accuracy of the transition from long-sleeve to short-sleeve, which is the purpose of our paper. Therefore, observation-based accuracy was added and evaluated.
4.2.3 Ablation study
Figure 10 shows the result of an ablation study analyzing the effect on the resulting quality of each major element comprising the CaP-VTON model. The same input image and reference clothing were applied in each configuration, and the short-sleeved silhou-ette and the degree of arm/shoulder skin exposure were compared in the output image. As a result, when using only Leffa, most of the existing long-sleeved clothing silhouettes re-mained present in the output image, and the exposure of the arms and shoulders was in-complete. In the case of Leffa + DressCode, the patterns and silhouettes of some reference clothing were reflected, but complete transformation was not made due to the influence of the existing long-sleeved structure. In the case of Leffa + Generate Skin, a significant part of the silhouette of the existing long-sleeved clothing remained, and the short-sleeved sil-houette was still incompletely reproduced. Finally, the CaP-VTON model of Leffa + DressCode + Generate Skin integrated all the components required to minimize the influ-ence of existing long-sleeved clothing, such that the short-sleeved silhouette and pattern of the reference clothing were naturally reflected.
Figure 10: Ablation study consisting of (a) Leffa only, (b) Leffa + DressCode, (c) Leffa + Generate Skin, and (d) CaP-VTON of Leffa + DressCode + Generate Skin. For each configuration, the same input image and reference clothing were applied to compare the short-sleeved silhouette and the degree of arm/shoulder skin exposure in the output image.
4.2.4 Evaluation based on Short-Sleeve Synthesis Probability
Leffa + DressCode and CaP-VTON were selected to quantitatively compare the results of short-sleeved clothing synthesis by the CaP-VTON model. The reasons for this are as follows. First, as pointed out at the beginning of this paper, Leffa’s detection is limited to bottoms and by the lack of connectivity to clothing types due to modified masking, so it is not suitable for use as a comparison target when evaluating the improvement induced by the proposed model. In other words, due to structural limitations, short-sleeved clothing conversion itself is not smoothly performed by this model, making it impossible to com-pare performance under the same conditions. Second, Leffa + Generate Skin is a configu-ration in which only a skin restoration module is applied without DressCode data, and it is insufficient to evaluate clothing silhouette conversion in the virtual try-on process.
@@ 4.2.4 Evaluation based on Short-Sleeve Synthesis Probability
Figure 11 is the result of comparing the output of the Leffa model and the CaP-VTON model learned with the DressCode dataset to quantitatively compare the short-sleeved clothing synthesis results of the proposed CaP-VTON model. In addition to the existing example images, real person images of the DeepFashion dataset were added to the image comparison to verify the quality of synthesis in more realistic situations.
The DeepFashion dataset is a dataset for the purpose of clothing recognition and at-tribute classification, and the purpose and characteristics of data are different from VI-TON-HD or DressCode configured for virtual try-on. Therefore, when applying Deep-Fashion data to a model that has already been sufficiently trained on VITON-HD or DressCode data, such as existing Leffa, the model tests the performance in new data and determines whether it can maintain stable performance in real person data or not as an objective evaluation index.
As can be seen from the figure, the silhouette consistency between short-sleeved and sleeveless clothing is more natural than that achieved by Leffa, and the boundary restora-tion of skin-exposed areas is smooth. In particular, when converting the long-sleeved in-put image to an output with short-sleeved clothing, it was observed that residual sleeves, parts of the body, or discontinuous textures remained on the arm in the existing Leffa, but in the proposed CaP-VTON model, the silhouette of the existing clothes was completely removed through the combination of Generate Skin pretreatment and multicategory masking, and the arm shape and skin tone were naturally restored.
In this study, CLIP-based clothing shape consistency evaluation was performed to automatically determine whether the reference short-sleeved clothing was actually syn-thesized into a normal short-sleeve image, with the intention of expanding to a long-sleeved input image. This evaluation was implemented by comparing the semantic similarity between images and texts using OpenAI’s CLIP (ViT-B/32) model. Specifically, each reference vector was generated by embedding two text prompts—”A person wearing a short sleeve” and ”A person wearing a long sleeve”. After that, for each composite image, the visual feature vector was extracted through the image encoder of CLIP, and the Cosine Similarity between the two text vectors was calculated. Through this approach, if the composite image showed a greater similarity to the “short sleeve” prompt, it was classi-fied as normal output, and on the contrary, if it showed a greater similarity to the “long sleeve” prompt, it was classified as abnormal output. In addition, if the difference between the two similarities (score diff = correct score - wrong score) was 0.02 or more, it was classified as normal with clear consistency, and if it was less than 0.02, it was partially classified as normal.
Figure 12: Comparison the percentage of correct output
Figure 12 visualizes the classification of the output results into normal output (when a short-sleeved silhouette consistent with the reference clothing is accurately created) and abnormal output (when a long-sleeved silhouette remains, or short-sleeved clothes are synthesized as long-sleeved clothes). After that, the ratios of each evaluation grade were aggregated to calculate the normal output ratio and abnormal output ratio for each model, and these were visualized in Figure 12.
@@
This result means that CAP-VTON recorded about a 15.4% greater clothing silhouette consistency accuracy than Leffa. In particular, CaP-VTON showed excellent performance in its ability to accurately reflect short-sleeved reference clothing after removing long sleeves and naturally restoring the skin of the arms and shoulders, which can be said to empirically prove the effectiveness of Generate Skin-based pre-processing and the im-proved masking pipeline.
@@
Meanwhile, Leffa showed excellent quantitative performance regarding the FID and SSIM, but the above experimental results suggest the limitation that it is difficult to fully explain the suitability of the model’s clothing replacement using only quantitative image quality evaluation. In practical applications, the accuracy and visual naturalness of the clothing shape transformation desired by the user are more important factors, and thus, the consistency-based evaluation method proposed in this study is a very important measure in a practical sense.
@@
In conclusion, CaP-VTON showed a marked improvement over Leffa regarding the structural consistency, meaning that the pipeline was designed to minimize reliance on existing clothing types and accurately reflect the style of reference clothing. These results support the feasibility of using CaP-VTON in high-dimensional applications such as full-body virtual wear and style transformation.